OpenStudy (anonymous):

What percentage is 99 of 714?

OpenStudy (anonymous):

99*714= ?

OpenStudy (anonymous):

70,686

OpenStudy (whpalmer4):

percentages are just proportions: \[\frac{99}{714} = \frac{x \%}{100\%}\]

OpenStudy (anonymous):

can yall help me with my question please

OpenStudy (anonymous):

I'm sorry I meant 99/714

OpenStudy (anonymous):

\[\frac{ 99 }{ 714}=0.138=13.8%\]

OpenStudy (anonymous):

0.13865546218

OpenStudy (anonymous):

how do you find percentage now?

OpenStudy (anonymous):

you go two decimal places to the right

OpenStudy (anonymous):

You divide the numbers?

OpenStudy (anonymous):

that gets you 13.9%

OpenStudy (anonymous):

@whpalmer4 I'm doing this correctly right?

OpenStudy (anonymous):

It's been a while... lol

OpenStudy (whpalmer4):

Yes, you can convert between decimals and percents by multiplying or dividing by 100. \[\frac{99}{714} = \frac{x}{100\%}\]Cross multiply to get \[99000\% = 714x\]Divide both sides by 714 \[x = 99000\%/714 = 13.8655\%\]

OpenStudy (anonymous):

awesome...

OpenStudy (whpalmer4):

remember that "percent" comes from "per centum" or "out of each 100". 3% means 3 out of each 100. if we don't have 100 (here we had 714), we just take the same ratio.

OpenStudy (whpalmer4):

714/100 = 7.14 so we had 7.14 "100s", and we divide 99 by that, 99/7.14 = 13.8655, so 13.8655 per 100. 13.8655%

OpenStudy (anonymous):

@faded911 make sense?

OpenStudy (anonymous):

yea